How might you determine the concentration of a strong acid using a standard potassium thiosulfate solution, potassium iodide, and a standardized solution of sodium thiosulfate?
1
Expert's answer
2013-01-08T11:58:23-0500
Thisis example of Redox Titration and connection between the reactantsis below
A standard reaction is as follows:
Acid and potassium iodide are added to a solution of potassium iodate getting the following reaction:
KIO3 + 5KI + 3H2SO4 = 3I2 + 3K2SO4 + 3H2O
represented by the following ionic equation:
IO3- + 5I- + 6H+ = 3I2 + 3H2O
Thiosulpathe is titrated against this solution (effectively against iodine):
I2 + 2Na2S2O3 = Na2S4O6 + 2NaI
represented by the following ionic equation:
I2 + 2S2O32- = S4O62- + 2I-
where the dark brown coloured solution of iodine turns pale yellow and finally colourless as the reaction proceeds (starch is used as indicator after the pale yellow transition forming a black solution due to an iodine-starch complex which turns colourless upon further addition of thiosulphate).